Sue bought two items for $8.10, one item cost $2.40 less than the other what did Sue pay for each item?

Answer by checkley79(3341)

Let x & x-$2.40 be the prices of the 2 items.
x+x-2.40=8.10
2x=8.10+2.40
2x=10.50
x=10.50/2
x=$5.25 for the most expensive item.
5.25-2.40=$2.85 for the least costly item.
Proof:
5.25+2.85=8.10
8.10=8.10